### v2

# SAVA: Scalable Learning-Agnostic Data Valuation

Abstract
Selecting data for training machine learning models is crucial since large, web-scraped, real datasets contain noisy artifacts that affect the quality and relevance of individual data points. These noisy artifacts will impact model performance. We formulate this problem as a data valuation task, assigning a value to data points in the training set according to how similar or dissimilar they are to a clean and curated validation set. Recently, LAVA  (just2023lava) demonstrated the use of optimal transport (OT) between a large noisy training dataset and a clean validation set, to value training data efficiently, without the dependency on model performance. However, the LAVA algorithm requires the entire dataset as an input, this limits its application to larger datasets. Inspired by the scalability of stochastic (gradient) approaches which carry out computations on batches of data points instead of the entire dataset, we analogously propose SAVA, a scalable variant of LAVA with its computation on batches of data points. Intuitively, SAVA follows the same scheme as LAVA which leverages the hierarchically defined OT for data valuation. However, while LAVA processes the whole dataset, SAVA divides the dataset into batches of data points, and carries out the OT problem computation on those batches. We perform extensive experiments, to demonstrate that SAVA can scale to large datasets with millions of data points and doesn’t trade off data valuation performance.
1 Introduction
Neural scaling laws empirically show that the generalization error decreases according to a power law as the data a model trains on increases. This has been shown for natural language processing, vision, and speech (kaplan2020scaling; henighan2020scaling; rosenfeld2019constructive; zhai2022scaling; radford2023robust). However, training neural networks on larger and larger datasets for moderate improvements in model accuracy is inefficient. Furthermore, production neural network models need to be continuously updated given new utterances that enter into common everyday parlance (lazaridou2021mind; baby2022incremental). It has been shown both in theory and in practice that sub-power law, exponential scaling of model performance with dataset size is possible by carefully selecting informative data and pruning uninformative data (sorscher2022beyond), and that generalization improves with training speed (lyle2020bayesian). Therefore, valuing and selecting data points that are informative: which have not been seen by the model, which do not have noisy labels, and which are relevant to the task we want to solve—are not outliers—can help to not only decrease training times, and reduce compute costs, but also improve overall test performance (mindermann2022prioritized; tirumala2023d4).
Figure 1: Overview of the proposed SAVA method. Instead of solving a single, but expensive, OT problem on the whole dataset, SAVA considers solving multiple cheaper smaller OT problems on batches of data points, and integrates the resulting calibrated gradients together for a data valuation task. $\text{OT}(\bar{\mu_{t}},\bar{\mu_{v}})$ denotes the OT cost and $\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})$ is its associated optimal transport plan. $\text{OT}(\mu_{B_{i}},\mu_{B_{j}^{\prime}})$ is the OT between a uniform distribution of points in the batch $B_{i}$ from the training set and the batch $B^{\prime}_{j}$ in the validation set. $s_{k}$ is SAVA’s final valuation score for training data point $z_{k}$. We provide a visualization of these artifacts generated by SAVA in LABEL:fig:sava_alg_artifacts.
Popular data selection and data pruning methods use variations of the model loss to value data points (jiang2019accelerating; pruthi2020estimating; paul2021deep). Crucially, these methods depend on the model used, and they are vulnerable to prioritizing data points with noisy labels or noisy features; data points that do not resemble the target validation set. When treating data valuation as a function of model performance, we introduce a dependency on a neural network model. This is the case when valuing a point using the leave-one-out (LOO) error, i.e., the change of model performance when the point is omitted from training. To rid our dependence on a neural network model, a promising idea is to leverage optimal transport (OT) between a training distribution and a clean validation distribution as a proxy to directly measure the value of data points in the training set in a model-agnostic fashion. In particular, the validation performance of each point in the training set can be estimated using the hierarchically defined Wasserstein between the training and the validation set (alvarez2020geometric; just2023lava). LAVA (just2023lava) has been shown to successfully value data by measuring the sensitivity of the hierarchically defined Wasserstein between training data points and validation distributions in a model-agnostic fashion. However, LAVA requires significant RAM consumption since its memory complexity grows quadratically $\mathcal{O}(N^{2})$ with the dataset size $N$. This hinders LAVA from scaling to large datasets.
In this paper, we present SAVA, a scalable variant of LAVA, for data valuation. Our method completely addresses the bottleneck of RAM requirements in LAVA. Intuitively, SAVA performs the OT computations on batches instead of on the entire dataset like LAVA (hence the analogy to the stochastic approach to (sub)gradient computation). Specifically, SAVA uses ideas from hierarchical optimal transport (yurochkin2019hierarchical; lee2019hierarchical) to enable OT calculations on batches instead of the entire dataset. We can scale up OT-based data valuation using SAVA to large real-world web-scrapped datasets. On benchmark problems SAVA performs comparably to LAVA while being able to scale to datasets two orders of magnitude larger without memory issues, while LAVA is limited due to hardware memory constraints.
Contributions:
1. We introduce a novel scalable data valuation method called SAVA that leverages the (sub)gradient of hierarchical optimal transport (yurochkin2019hierarchical; lee2019hierarchical) which performs OT computations on batches of data points, enabling OT-based data valuation to large datasets.
2. We correct the essential theoretical result on the trade-off for using entropic regularization to compute the OT (sub)gradient in LAVA (just2023lava, Theorem 2).111see Appendix LABEL:appendix_issue_theorem2_lava for the discussion Consequently, by building upon our refined theory, we derive the exact trade-off to estimate the (sub)gradient of hierarchical OT with entropic regularization in our framework.
3. We provide an extensive experimental analysis to demonstrate the improved scalability with increasing dataset sizes with respect to baselines.
2 Optimal Transport for Data Valuation
2.1 Optimal Transport for Labeled Datasets
Let ${\mathcal{X}}$ be the feature space, and $V$ be the number of labels. We write $f_{t}:{\mathcal{X}}\mapsto\left\{0,1\right\}^{V}$ and $f_{v}:{\mathcal{X}}\mapsto\left\{0,1\right\}^{V}$ for the labeling functions for training and validation data respectively. Given the training set ${\mathbb{D}}_{t}=\left\{(x_{i},f_{t}(x_{i}))\right\}_{i=1}^{N}$ and the validation set ${\mathbb{D}}_{v}=\left\{(x^{\prime}_{i},f_{v}(x^{\prime}_{i}))\right\}_{i=1}^{N^{\prime}}$, the corresponding measures for sets ${\mathbb{D}}_{t},{\mathbb{D}}_{v}$ are $\mu_{t}(x,y)=\frac{1}{N}\sum_{i=1}^{N}\delta_{(x_{i},y_{i})}$ and $\mu_{v}(x^{\prime},y^{\prime})=\frac{1}{N^{\prime}}\sum_{i=1}^{N^{\prime}}\delta_{(x^{\prime}_{i},y^{\prime}_{i})}$ respectively where $\delta$ is the Dirac function, and $y,y^{\prime}$ are labels of $x,x^{\prime}$ respectively. For simplicity, let ${\mathcal{Z}}=({\mathcal{X}},{\mathcal{Y}})$ where ${\mathcal{Y}}$ is the space of labels. For ease of reading, we summarize all notations in Appendix Table LABEL:table:notation.
Following alvarez2020geometric, we compute the distance between two labels by leveraging the OT distance between the conditional distributions of the features given each label, i.e., $\mu_{t}(\cdot|y_{t})=\frac{\mu_{t}(\cdot)I[f_{t}(\cdot)=y_{t}]}{\int\mu_{t}(\cdot)I[f_{t}(\cdot)=y_{t}]}$ for label $y_{t}$ in $\mu_{t}$, where $I$ is the indicator function.
Let d be the metric of the feature space ${\mathcal{X}}$, e.g., the Euclidean distance. The distance between labels $y_{t}$ and $y_{v}$ is $\textnormal{OT}_{{\textnormal{d}}}(\mu_{t}(\cdot|y_{t}),\mu_{v}(\cdot|y_{v}))$, i.e., the metric of the label space ${\mathcal{Y}}$. Consequently, the cost between feature-label pairs in ${\mathcal{Z}}=({\mathcal{X}},{\mathcal{Y}})$ is ${\textnormal{C}}((x_{t},y_{t}),(x_{v},y_{v}))={\textnormal{d}}(x_{t},x_{v})+c\,\textnormal{OT}_{{\textnormal{d}}}(\mu_{t}(\cdot|y_{t}),\mu_{v}(\cdot|y_{v}))$, where $c>0$ is a weight coefficient. Therefore, with the cost matrix C, we can use the OT on the represented measures $\mu_{t},\mu_{v}$ to compute the distance between the training and validation sets, i.e., $d({\mathbb{D}}_{t},{\mathbb{D}}_{v})$, without relying on external models or parameters as follows
[…]
where $\Pi(\mu_{t},\mu_{v})$ is the set of transportation couplings with marginals as $\mu_{t}$ and $\mu_{v}$. To simplify notations, we drop C, and use OT when the context is clear. We further write $\pi^{*}$ for the optimal transport plan in Eq. (1). In practice, we leverage the entropic regularization (Cuturi-2013-Sinkhorn) to reduce the OT complexity into quadratic (from super cubic) w.r.t. the number of input supports, defined as
[…]
Additionally, the OT problem in Eq. (1) is a constrained convex minimization, it is naturally paired with a dual problem, i.e., constrained concave maximization problem, as follows:
$\textnormal{OT}_{{\textnormal{C}}}(\mu_{t},\mu_{v})=\max_{(f,g)\in{\mathcal{R}}({\textnormal{C}})}\left<f,\mu_{t}\right>+\left<g,\mu_{v}\right>,$ (3)
[…]
2.2 LAVA: Data valuation via calibrated OT gradients
As mentioned in just2023lava, the OT distance is known to be insensitive to small differences while also being not robust to large deviations. This feature is naturally suitable for detecting abnormal data points, i.e., disregarding normal variations in distances between clean data while being sensitive to abnormal distances of outlying points. Therefore, the (sub)gradient of the OT distance w.r.t. the probability mass associated with each point can be leveraged as a surrogate to measure the contribution of that point. More precisely, the (sub)gradient of the OT distance w.r.t. the probability mass of data points in the two datasets can be expressed:
$\nabla_{\mu_{t}}\textnormal{OT}(f^{*},g^{*})=(f^{*})^{T},\hskip 3.99994pt\nabla_{\mu_{v}}\textnormal{OT}(f^{*},g^{*})=(g^{*})^{T}.$ (4)
The dual solution is unique, up to a constant due to the redundant constraint $\sum_{i=1}^{N}\mu_{t}(z_{i})=\sum_{i=1}^{M}\mu_{v}(z^{\prime}_{i})=1$. Therefore, for measuring the subgradients of the OT w.r.t. the probability mass of a given data point in each dataset, just2023lava we calculate the calibrated gradients (i.e., a sum of all elements equals to a zero vector) [^2] To remove the degree of freedom which comes from the fact that one among all row/column sum constraints for the transport polytope is redundant, among the OT subgradients, we fix the zero-sum subgradient following the convention in Cuturi-2014-Fast and the calibrated approach in LAVA (just2023lava). as
$\frac{\partial\textnormal{OT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{i})}=f^{*}_{i}-\sum_{j\in\left\{1,...,N\right\}\textbackslash i}\frac{f^{*}_{j}}{N-1}.$ (5)
The calibrated gradients predict how the OT distance changes as more probability mass is shifted to a given data point. This can be interpreted as a measurement of the contribution of the data point to the OT. Additionally, if we want a training set to match the distribution of the validation dataset, then either removing or reducing the mass of data points with large positive gradients, while increasing the mass of data points with large negative gradients can be expected to reduce their OT distance. Therefore, as demonstrated in just2023lava, the calibrated gradients provide a powerful tool to detect and prune abnormal or irrelevant data in various applications.
Memory Limitation. While being used with the current best practice, the Sinkhorn algorithm for entropic regularized OT (Cuturi-2013-Sinkhorn) still runs in quadratic memory complexity $O(N^{2})$ with the dataset size $N$, as it requires performing operations on the entire dataset, using the full pairwise cost matrix. Consequently, the memory and RAM requirements for the Sinkhorn algorithm primarily depend on the dataset size $N$. Additionally, notice that a dense square (float) matrix of size $N=10^{5}$ will require at least $74$ GB of RAM and $N=10^{6}$ will take $7450$ GB of RAM which is prohibitively expensive. Therefore, LAVA is limited to small datasets.
Scalability. Inspired by the scalability of stochastic (gradient) approaches where the computation is carried out on batches of data points instead of the whole dataset as in the traditional gradient, we follow this simple but effective scheme to propose an analog for OT, named SAVA which is a scalable variant of LAVA with its OT computation on batches. Intuitively, SAVA also leverages the hierarchically defined OT as in LAVA, but it performs OT on batches of data points instead of on the entire dataset as in LAVA. We introduce SAVA in the next section.
The scalable OT-based data valuation we introduced in this section focuses on the (sub)gradient of the OT instead of the OT distance itself. Therefore, popular scalable OT approaches, sliced-Wasserstein (rabin2011wasserstein) or Sobolev transport (le2022st), may not be suitable. In the next section, we introduce our novel scalable approach for computing the (sub)gradient of the OT using hierarchical OT (yurochkin2019hierarchical; lee2019hierarchical). We focus on the problem of data valuation but our work can be applied to other large dataset applications where the (sub)gradient of the OT is required.
3 SAVA: Scalable Data Valuation
We present SAVA, a scalable data valuation method, scaling LAVA to large-scale datasets. Instead of solving a single (but expensive) OT problem for distributions on the entire datasets, i.e., $\textnormal{OT}(\mu_{t},\mu_{v})$ in LAVA with the pairwise cost matrix size ${\mathbb{R}}^{N\times N^{\prime}}$, we consider solving multiple (cheaper) OT problems for distributions on batches of data points. For this purpose, our algorithm performs data valuation on two levels of hierarchy: across batches, and across data points within two batches. Thus, SAVA can complement LAVA for large-scale applications.
Algorithm 1 Scalable Data Valuation (SAVA)
Input: a threshold $\tau$, $\varepsilon$ for Sinkhorn algorithm, let $z=(x,y)$
Output: training data values $s_{k}$ for all $k\in[N_{i}]$ for all $i\in[K_{t}]$.
for $i=1,...,K_{t}$ do
      2 for $j=1,...,K_{v}$ do
6            543 Compute ${\textnormal{C}}_{kl}\bigl{(}B_{i},B^{\prime}_{j}\bigr{)},\forall k\in[N_{i}],\forall l\in[N^{\prime}_{j}]$ by using Eq. (7). Compute ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})},g^{*}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$ by solving $\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$. Set ${\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}=\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$.           // distance $d({B}_{i},{B}^{{}^{\prime}}_{j})$ on batches.
7
1098Compute ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}\in{\mathbb{R}}^{K_{t}\times K_{v}}$ by solving $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ for $i=1,...,K_{t}$ do
12      11 for $k=1,...,N_{i}$ do
15            1413 Compute $\frac{\partial\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})},\,\,\forall j\in[K_{v}]$ using Eq. (13). Compute $s_{k}=\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{l})}$ using Eq. (12). // valuation score for $z_{k}\in B_{i}$.
16
Hierarchical OT. We follow the idea in hierarchical OT approach (yurochkin2019hierarchical; lee2019hierarchical) to partition the training dataset ${\mathbb{D}}_{t}$ of $N$ samples into $K_{t}$ disjoint batches ${B}=\left\{{B}_{i}\right\}_{i=1}^{K_{t}}$. Similarly, for the validation set ${\mathbb{D}}_{v}$ of $N^{\prime}$ samples into $K_{v}$ disjoint batches ${B}^{\prime}=\left\{{B}^{\prime}_{j}\right\}_{j=1}^{K_{v}}$. Additionally, for all $i\in[K_{t}],j\in[K_{v}]$, let the number of samples in batches $B_{i},B^{\prime}_{j}$ as $N_{i},N^{\prime}_{j}$ respectively. The corresponding measures of the training and validation sets w.r.t. the batches are defined as: $\bar{\mu}_{t}({B})=\frac{1}{K_{t}}\sum_{i=1}^{K_{t}}\delta_{({B}_{i})}$ and $\bar{\mu}_{v}({B}^{\prime})=\frac{1}{K_{v}}\sum_{j=1}^{K_{v}}\delta_{({B}^{\prime}_{j})}$ respectively. We then define a distance between the datasets as the hierarchical optimal transport (HOT) between the measures $\mu_{t},\mu_{v}$ as OT distance for corresponding represented measures on batches, i.e., $\bar{\mu}_{t}$ and $\bar{\mu}_{v}$ as in Sec. 2.1 as follows:
$d(\mu_{t},\mu_{v}):=\textnormal{HOT}(\mu_{t},\mu_{v}):=\textnormal{OT}(\bar{\mu}_{t},\bar{\mu}_{v}).$ (6)
It is worth noting that HOT finds the optimal coupling at the batch level, but not at the support data point level as in the classic OT. Therefore, it can be seen that
$\text{OT}(\mu_{t},\mu_{v})\leq\text{HOT}(\mu_{t},\mu_{v}).$
The equality happens when either each batch only has one support or each dataset only has one batch.
Our goal is to estimate the (sub)gradient $\frac{\partial d(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}=\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}$ where $\textnormal{HOT}(\mu_{t},\mu_{v})$ is defined in the Eq. (6). Computing this derivative involves two OT estimation steps including (i) OT between individual data points within two batches to compute $d({B}_{i},{B}^{\prime}_{j}):=\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$, where $\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}}$ are corresponding measures for batches ${B}_{i},{B}^{\prime}_{j}$ respectively, and subsequently a cost matrix ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$ on pairwise batches for input measures over batches; (ii) $d(\mu_{t},\mu_{v})=\textnormal{HOT}(\mu_{t},\mu_{v}):=\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$.
Pairwise cost between batches. We estimate the distance between two batches as the OT problem between ${B}_{i}$ and ${B}_{j}$, i.e., $d({B}_{i},{B}^{\prime}_{j}):=\textnormal{OT}({B}_{i},{B}^{\prime}_{j})$ as discussed in Sec. 2.1 by viewing the OT problem between two labeled (sub)datasets or batches ${B}_{i}$ and ${B}^{\prime}_{j}$.
More precisely, to solve this OT problem, we calculate the pairwise cost for data points between two batches ${\textnormal{C}}_{kl}(B_{i},B^{\prime}_{j})\in{\mathbb{R}}^{N_{i}\times N^{\prime}_{j}}$, where $\forall k\in[N_{i}],\forall l\in[N^{\prime}_{j}]$, the element ${\textnormal{C}}_{kl}(B_{i},B^{\prime}_{j})$ is the cost between two labeled data points $(x_{k},y_{k})\in B_{i}$ and $(x^{\prime}_{l},y^{\prime}_{l})\in B^{\prime}_{j}$, calculated as
${\textnormal{C}}_{kl}\bigl{(}B_{i},B^{\prime}_{j}\bigr{)}={\textnormal{d}}(x_{k},x^{\prime}_{l})+c\,\textnormal{OT}_{{\textnormal{d}}}\bigl{(}\mu_{B_{i}}(\cdot|y_{k}),\mu_{B^{\prime}_{j}}(\cdot|y^{\prime}_{l})\bigr{)}.$ (7)
Given the cost matrix ${\textnormal{C}}(B_{i},B^{\prime}_{j})$, we solve $\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{{}^{\prime}}_{j}})$ to get dual solutions ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})},g^{*}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$, and the OT distance for $d({B}_{i},{B}^{\prime}_{j})$, i.e., $\textnormal{OT}_{{\textnormal{C}}}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$.
We repeat this process and solve the OT problem for each pair $(B_{i},B^{\prime}_{j})$, i.e., the OT problem for distributions on batches of data points, across the training and validation datasets, for all $i\in[K_{t}],j\in[K_{v}]$. This enables us to define the cost matrix for pairwise batches in $\bar{\mu}_{t},\bar{\mu}_{v}$, denoted as ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}(\bar{\mu}_{t},\bar{\mu}_{v})\in{\mathbb{R}}^{K_{t}\times K_{v}}_{+}$ where we recall that $K_{t}$ and $K_{v}$ are the number of batches in training and validation sets. Hence, for this cost matrix ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$, the element ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}_{ij}$ is computed as $\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$, for all $i\in[K_{t}],\forall j\in[K_{v}]$.
Batch valuation. Given the pairwise cost matrix across batches ${\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}$, we compute the data valuation for each batch via the (sub)gradient of the distance $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ w.r.t. the probability mass of batches in the two datasets, i.e., $\frac{\partial\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\partial\bar{\mu}_{t}({B}_{i})}$. These partial derivatives measure the contribution of the batches to the OT distance, i.e., shifting more probability mass to the batch would result in an increase or decrease of the dataset distance, respectively.
More precisely, let $\bar{f}^{*},\bar{g}^{*}$ be the optimal dual variables of $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$, then the data valuation of the batch $B_{i}$ in the training set ${\mathbb{D}}_{t}$ is estimated as follows:
$\frac{\partial\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\partial\bar{\mu}_{t}({B}_{i})}=\bar{f}_{i}^{*}.$ (8)
Since the optimal dual variables are only unique up to a constant, we follow just2023lava to normalize these optimal dual variables such that the sum of all elements is equal to a zero vector:
$\frac{\partial\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\partial\bar{\mu}_{t}(B_{i})}=\bar{f}_{i}^{*}-\sum_{j\in\left\{1,...,K_{t}\right\}\backslash i}\frac{\bar{f}^{*}_{j}}{K_{t}-1}.$ (9)
Using batch valuation for data point valuation. After solving the OT problem over batches, we obtain the optimal transport plan ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ which is used to compute the data valuation over individual data points:
${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}=\text{diag}{(\bar{u}^{*})}\exp\Bigl{(}-\frac{{\color[rgb]{.5,0,.5}\bar{\textnormal{C}}}(\bar{\mu}_{t},\bar{\mu}_{v})}{\varepsilon}\Bigr{)}\text{diag}{(\bar{v}^{*})},$ (10)
where $\text{diag}(\cdot)$ is matrix diagonal operator, $(\bar{u}^{*},\bar{v}^{*})$ is a dual solution from Sinkhorn algorithm. Additionally, we have $\bar{u}^{*}=\exp(-\frac{1}{2}-\frac{\bar{f}^{*}}{\varepsilon}),\bar{v}^{*}=\exp(-\frac{1}{2}-\frac{\bar{g}^{*}}{\varepsilon})$ following Cuturi-2013-Sinkhorn.
For a data point $z\in{B}_{i}\subset{\mathbb{D}}_{t}$, its data valuation score can be computed as follows:
$\sum_{j=1}^{K_{v}}{\color[rgb]{1,.5,0}\bar{\pi}^{*}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}\frac{\partial\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z)}.$ (11)
Using HOT, we can measure the (sub)gradients of the OT  distance w.r.t. the probability mass of a given data point in each dataset via the calibrated gradient summarized in the following Lemma 1. We refer to Appendix Table LABEL:table:notation for the notations.
Lemma 1.
The calibrated gradient for a data point $z_{k}$ in the batch ${B}_{i}$ in ${\mathbb{D}}_{t}$ can be computed as:
$\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}=\sum_{j=1}^{K_{v}}{\color[rgb]{1,.5,0}\bar{\pi}^{*}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}\frac{\partial\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})},$ (12)
where the calibrated gradient of OT for measures on batches is calculated as follows:
$\frac{\partial\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})}={\color[rgb]{1,0,0}f^{*}_{k}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}\hskip 10.00002pt-\sum_{l\in\left\{1,...,N_{i}\right\}\setminus k}\frac{{\color[rgb]{1,0,0}f^{*}_{l}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}}{N_{i}-1},\forall j\in[K_{v}].$ (13)
It is common practice to compute the OT via its entropic regularization, using the Sinkhorn algorithm (Cuturi-2013-Sinkhorn). We quantify the deviation in the calibrated gradients caused by the entropy regularizer. This analysis shows the potential impact of the deviation on applications built on these gradients.
We refine the theoretical result (just2023lava, Theorem 2), which uses entropic regularization to approximate the OT (sub)gradient in two aspects. Firstly, we take into account the non-negative constraint of the OT plan $\pi\geq 0$ in the dual formulation. Secondly, we correct the discrete formulation for entropic regularization (see LABEL:appendix_issue_theorem2_lava for the discussion).
Theorem 2 (Refined Theorem 2 in just2023lava).
The difference between calibrated gradients for two data points in ${\mathbb{D}}_{t}$ and ${\mathbb{D}}_{v}$ can be calculated as
$\frac{\partial\operatorname{OT}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{i}\right)}-\frac{\partial\operatorname{OT}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{k}\right)}=\frac{\partial\textnormal{OT}_{\varepsilon}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{i}\right)}-\frac{\partial\textnormal{OT}_{\varepsilon}\left(\mu_{t},\mu_{v}\right)}{\partial\mu_{t}\left(z_{k}\right)}$
$\hskip 170.00026pt+\varepsilon\frac{N}{N-1}\left({\color[rgb]{0,0,1}\log\frac{\left(\pi_{\varepsilon}^{*}\right)_{ij}/\mu_{t}(z_{i})}{\left(\pi_{\varepsilon}^{*}\right)_{kj}/\mu_{t}(z_{k})}+\frac{h_{kj}^{*}-h_{ij}^{*}}{\varepsilon}}\right),$ (14)
where $h^{*}$ is the corresponding optimal dual variable, accounting for non-negative constraint on the transportation plan $\pi\geq 0$ in the primal OT formulation.
Proof.
Refer to Appendix LABEL:appendix_proof_theorem2 for the proof. ∎
The term in the brackets (in blue) is the key difference between our new result and the previous one derived in LAVA (just2023lava, Theorem 2).
Lemma 3.
Let $\textnormal{HOT}(\mu_{t},\mu_{v})$ and $\textnormal{OT}_{\varepsilon}(\mu_{t},\mu_{v})$ be the HOT and entropic regularized OT between the measures $\mu_{t}$ and $\mu_{v}$ associated with the two datasets ${\mathbb{D}}_{t}$ and ${\mathbb{D}}_{v}$ respectively. The difference between the calibrated gradients for two data points $\{z_{l},z_{h}\}\in B_{i}\subset{\mathbb{D}}_{t}$ can be calculated as
$\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{k})}-\frac{\partial\textnormal{HOT}(\mu_{t},\mu_{v})}{\partial\mu_{t}(z_{l})}=\sum_{j=1}^{K_{v}}{\color[rgb]{1,.5,0}\bar{\pi}^{*}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}\Bigl{[}\frac{\partial\textnormal{OT}_{\varepsilon}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{k})}-\frac{\partial\textnormal{OT}_{\varepsilon}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}{\partial\mu_{{B}_{i}}(z_{l})}$
$\hskip 150.00023pt+\varepsilon\frac{N_{i}}{N_{i}-1}\left(\log\frac{(\bar{\pi}^{*}_{\varepsilon})_{k,j}/\mu_{t}(z_{k})}{(\bar{\pi}^{*}_{\varepsilon})_{l,j}/\mu_{t}(z_{l})}+{\frac{h_{lj}^{*}-h_{kj}^{*}}{\varepsilon}}\right)\Bigr{]}.$ (15)
Proof.
Refer to Appendix LABEL:appendix_proof_lemma3 for the proof. ∎
We make a similar observation in LAVA that the ground-truth (sub)gradient difference between two training points $z_{k}$ and $z_{l}$ is calculated based on the HOT formulation and can be approximated by the entropic regularized formulation $\textnormal{OT}_{\varepsilon}$, such as via the Sinkhorn algorithm (Cuturi-2013-Sinkhorn). In other words, we can calculate the ground-truth difference based on the solutions to the regularized problem plus some calibration terms that scale with $\varepsilon$. In addition, in our case with HOT, the (sub)gradient difference also depends on the additional optimal assignment across batches ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ which is again estimated by the Sinkhorn algorithm.
4 Properties and Discussions
The SAVA Algorithm. We outline the computational steps of SAVA in Algorithm 1. From Lines 1 to 5, we solve multiple OT tasks for data points between two batches $B_{i},B^{\prime}_{j}$. We obtain the dual solution ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}\in{\mathbb{R}}^{N_{i}}$. Additionally, these OT distances are used to fill in the cost matrix for pairwise batches ${\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}_{ij}(\bar{\mu}_{t},\bar{\mu}_{v})}=\textnormal{OT}(\mu_{B_{i}},\mu_{B^{\prime}_{j}})$. Here, the cost matrix C is of size $N_{i}\times N^{\prime}_{j}$ where the batch sizes $N_{i},N^{\prime}_{j}\ll N$. The required memory complexity is at least $\tilde{\mathcal{O}}(N^{2}_{i})$.
In line 6 Algorithm 1, we solve the OT problem across batches $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ to obtain ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$. The cost matrix is small, of size $K_{t}\times K_{v}$, so less expensive compared to working with the full dataset just2023lava. Finally, from Lines 7 to 10, we estimate valuation scores for training data using the auxiliary matrices computed in the previous steps, including ${\color[rgb]{1,0,0}f^{*}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})}$ and ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$.
SAVA Memory Requirements. The Sinkhorn algorithm computational complexity is at least $\tilde{\mathcal{O}}(N^{2})$ where $N$ is the number of samples. This comes from the main OT  step of $\textnormal{OT}(\mu_{t},\mu_{v})$ over the full cost matrix C of size $N\times N^{\prime}$ to subsequently calculate the calibrated gradient in Eq. (5) where $N$ and $N^{\prime}$ are the training and validation dataset sizes. SAVA overcomes this limitation by solving multiple smaller OT problems $\textnormal{OT}(\mu_{{B}_{i}},\mu_{{B}^{\prime}_{j}})$ on distributions for batches of data points only.
Practical Implementation with Caching. While yielding a significant memory improvement, SAVA’s runtime is not necessarily faster than LAVA. To speed up SAVA, we propose to implement Algorithm 1 by caching the label-to-label costs between points in the validation and training batches: $\textnormal{OT}_{{\textnormal{d}}}(\cdot,\cdot)$ in Sec. 2.1 so that it is only calculated once in the first iteration of Algorithm 1 and reused for subsequent batches. This significantly reduces SAVA runtimes with no detriment to performance LABEL:fig:l2l_caching_ablation. All experimental results in Section 5, unless otherwise stated, implement this caching strategy.
Batch Sizes. If we consider Eq. (6), HOT provides the upper bound for the OT since its optimal coupling is on batches of data points. HOT recovers the OT when either batch only has one support or each dataset only has one batch. Consequently, up to a certain batch size, when increasing the batch size $N_{i}$, HOT converges to the OT, but its memory complexity also increases, i.e., $\tilde{\mathcal{O}}(N^{2}_{i})\rightarrow\tilde{\mathcal{O}}(N^{2})$. On the other hand, if the batch size $N_{i}$ is too small, the number of batches $K_{t}$ will be large. As a result, the memory complexity will also be high (i.e., for solving $\textnormal{OT}_{{\color[rgb]{.5,0,.5}\bar{{\textnormal{C}}}}}(\bar{\mu}_{t},\bar{\mu}_{v})$ in Algorithm 1). Thus, the batch size will trade off the memory complexity and the approximation of HOT to OT.
5 Experiments
We aim to the test two following hypotheses: 1. Can SAVA scale and overcome the memory complexity issues which hinder LAVA while maintaining similar performance? 2. Can SAVA scale to a large real-world noisy dataset with over a million data points?
[…]
5.1 Dataset Corruption Detection
We test the scalability of SAVA versus LAVA (just2023lava) by leveraging the CIFAR10 dataset, introducing a corruption to a percentage of the training data, but keeping the validation set clean. We then assign a value to each data point in the training set. Following pruthi2020estimating; just2023lava, we sort the training examples by their values. An effective data valuation method would rank corrupted examples in a prefix of ordered points. We use the fraction of corrupted data recovered by the prefix of size $N/4$ as our detection rate (see an example in LABEL:app:data_val_rankingsLABEL:fig:data_val_example).
Setup. We consider $4$ different corruptions (see LABEL:appendix_data_corruptions for details) applied to training data following just2023lava: (i) noisy labels, (ii) noisy features, (iii) backdoor attacks and (iv) poison detections. All experiments are run on a Tesla K80808080 Nvidia GPUs with $12$GB GPU RAM. Unless otherwise stated, all results reported are a mean ±$\pm 1$ standard deviation over $5$ independent runs.
Baselines. Our main baseline is LAVA. For SAVA and LAVA we use features from a pre-trained ResNet18 (he2016deep). For SAVA, we use a default batch size of $N_{i}=1024$ which is its main hyperparameter. [^3] See LABEL:sec:ablations for the study of the sensitivity of the batch size $N_{i}$ in SAVA. We consider KNN Shapley (jia2019efficient), the Shapley value measures the marginal improvement in the utility of a data point and uses KNN to approximate the Shapley value. KNN Shapley also uses ResNet18 features to calculate Shapley values, we tune $k$ as its performance is sensitive to this hyperparameter. We consider TracInCP (pruthi2020estimating) which measures the influence of each training point by measuring the difference in the loss at the beginning versus the end of training. We also compare with Influence Function (koh2017understanding) which approximates the effect of holding out a training point on the test error, it uses expensive approximations of the Hessian to calculate the influence of a training point. We also consider Data-OOB (kwon2023data) which uses bagging estimators to value data points, data points are embedded using a pre-trained feature extractor, and the bagging estimator is a decision tree (see LABEL:sec:imp_details_cifar10 for implementation details). We finally consider a naive OT baseline which obtains values for data points at a batch level, and aggregates values across validation batches; the baseline essentially performs LAVA only at a data point level within a batch and averaging values across validation batches. We call this baseline Batch-wise LAVA. Although Batch-wise LAVA obtains good corruption detection results, it is very sensitive to the batch size (see LABEL:sec:batchwise_lava_robustness).
Noisy labels. We corrupt $30\%$ of the labels in the training set by randomly assigning the target a different label. From Figure 2, we can see that LAVA has an out-of-memory (OOM) issue when valuing the full training set of $50$K points. In contrast, SAVA, Batch-wise LAVA, KNN Shapley and Data-OOB can consistently value and detect all corruptions when inspecting $12.5$K ordered samples by values. Influence Functions matches the performance of SAVA, but is very expensive to run on large datasets. TracInCP is unable to detect noisy labels better than random selection, similar to the observations in just2023lava. Consequently, when pruning $30\%$ of the data, SAVA can consistently improve its accuracy on the test set as the training set size increases LABEL:sec:cifar10_pruning.
Noisy features. We add Gaussian noise to $30\%$ of the training images to simulate feature corruptions that might occur in real datasets. From Figure 2, LAVA obtains good performance for moderate dataset sizes, but for training set sizes above $10$K, some runs have OOM errors when embedding the entire training set into memory and when calculating the full cost matrix for OT problem. In contrast, SAVA can consistently value and detect corruptions. Batch-wise LAVA is also able to scale and gets similar performance to SAVA. As can KNN Shapley and Data-OOB, albeit its detection rate is lower than SAVA; as we prune $30\%$ of the data for larger and larger dataset sizes performance of SAVA outperforms KNN Shapley LABEL:fig:cifar10_corruptions_prune. TracInCP and Influence Functions both struggle to detect noisy data points, both methods were originally shown to detect noisy labels, so we do not expect them to work well beyond noisy label detection.
Backdoor attacks. We corrupt $10\%$ of the data with a Trojan square attack (liu2018trojaning). SAVA, Batch-wise LAVA and KNN Shapley can scale as the dataset size increases while LAVA has an OOM error when valuing the largest dataset with $50$k points. TracInCP and Influence Functions both struggle to detect backdoored training points and have long runtimes. Data-OOB also struggles to value data points in this scenario, this is in line with recent surveys which show less strong performance when valuing data with noisy features jiang2023opendataval.
Poisoning attacks. We corrupt $10\%$ of the data with a poison frogs attack (shafahi2018poison). We find in Figure 2 that SAVA and Batch-wise LAVA can scale and maintain a high detection rate. In contrast, KNN Shapley struggles to detect corrupted data points after inspecting $25\%$ of the ordered training data. TracInCP, Data-OOB and Influence Functions struggle to detect the corrupted data.
5.2 Large Scale Valuation and Pruning
We test our second hypothesis: whether SAVA can scale to a large real-world dataset. We consider the web-scrapped dataset Clothing1M (xiao2015learning) where the training set has over $1$M images whose labels are noisy and unreliable. However, the validation set has been curated. Clothing1M is a $14$-way image classification problem and has been used for previous work on online data selection (mindermann2022prioritized). We consider SAVA and other data pruning methods as baselines to remove low-value training data points before training a ResNet18 classifier.
Figure 3: SAVA can scale to a large web-scrapped dataset. We use SAVA and other baselines, to value data points and then prune a certain percentage of the noisy training set. The resulting dataset is used for training a classifier.
We compare to Batch-wise LAVA, introduced in Section 5.1. We also compare to EL2N (paul2021deep) which values training points using the loss on several partially trained networks to decide which points to remove. We train $10$ models for $10$ epochs, we perform a cross-validate a sliding window of values to decide which EL2N values to keep (see LABEL:sec:imp_details_el2n for details). We also consider supervised prototypes (sorscher2022beyond) which prunes image embeddings according to how similar they look to cluster centers after clustering image embeddings (see LABEL:sec:imp_details_sup_proto for details).
If we were to train a classifier on the full noisy training set, we would obtain an accuracy of $67.6\pm 0.2$, this remains constant as we randomly prune more data. Supervised prototypes obtains better results than random, and improves slightly over random pruning. This is expected since supervised prototypes is a semantic deduplication method and ignores label information. SAVA, Batch-wise LAVA and EL2N perform well and we find that SAVA performs better than both EL2N and Batch-wise LAVA. This shows the benefit of the SAVA’s weighting of the (sub)gradient of the OT across the validation dataset using optimal plan across batches ${\color[rgb]{1,.5,0}\bar{\pi}^{*}(\bar{\mu}_{t},\bar{\mu}_{v})}$ rather than Batch-wise LAVA’s uniform weighting. SAVA produces the best accuracy model of $70.0\pm 0.2$ in Figure 3.
6 Conclusions
We have presented a scalable extension to LAVA to address the challenges posed by large-scale datasets. Instead of relying on the expensive OT computation on the whole dataset, our proposed SAVA algorithm involves jointly optimizing multiple smaller OT tasks across batches of data points and within individual batches. We empirically show that SAVA maintains performance in data valuation tasks while successfully scaling up to handle a large real-world noisy dataset. This makes the data valuation task feasible for large-scale datasets. Our proposed method, along with the theoretical correction (just2023lava, Theorem 2), will complement, strengthen and improve (in terms of efficiency) OT-based data valuation.
Limitations. HOT finds the optimal coupling at the batch level, but not at the global level as in the traditional OT, i.e., $\text{OT}(\mu_{t},\mu_{v})\leq\text{HOT}(\mu_{t},\mu_{v})$, which makes the validation error bound looser (just2023lava, Eq. (1), Theorem 1). However, our experiments indicate little performance degradation for small datasets between SAVA and LAVA. See LABEL:app:limitations for further discussion on limitations.
Reproducibility Statement
We have described in detail our algorithm in Section 3 with Algorithm 1. All implementation details for our method and baselines are described in Section 5 and expanded on in LABEL:sec:impl_details. The source code to reproduce our experiments is included in our submission supplementary material. We will open-source our code upon acceptance of our paper.